One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Write the letter of the larger number.
a.17
b.5
c.34
d.7

Answer by junior403(76)

First...
One number (n) is 12 more than another (n+12).
The sum of the smaller number (n)
and twice the larger number + 2(n+12)
is 39 (=39)
So lets put that formula together.
n + 2(n + 12) = 39
Distribute...
n + 2n + 24 = 39
Subtract 24 from each side...
n + 2n = 15
Add like terms...
3n = 15
Divide both sides by 3...
n = 5
So, the smaller number (n) = 5
The larger number (n + 12) = 17
The correct answer is... a. 17